# Forms of Understanding for XAI-Explanations

### v2

@@ 3.1 Shallow comprehension, shallow enabledness
We see the combination of shallow comprehension and shallow enabledness as a potential starting point for understanding. Thus, this section will focus on the mental state before a change in understanding, or here, before explanation. The first step is to define the underlying conditions and premises of this stage. Any process of understanding should begin with and be tailored to the explainee’s mental representation model, which is built from (prior) own and mediated experiences with the domain and prior knowledge about familiar things. At the beginning of an explanation, explainers and explainees need to establish ’common ground’ [25, p. 109], which involves assessing “what the other person knows at the start, then tailoring one’s own contributions to add to the relevant common ground for that occasion […]”. Common ground is usually based on the prior knowledge that explainees bring to the explanation process. Prior knowledge thus influences the course of the explanation and the understanding that follows: Explainers can build their explanation on it, and explainees’ mental processing of new knowledge is influenced by it [22]. Therefore, it is particularly interesting to know what prior knowledge is and which parts of it are relevant for an explanation, and how the explainer can identify these relevant parts of prior knowledge. To clarify the situation in this field, we begin with the two examples.
@@
We can imagine that a novice in the field of shallow comprehension and shallow enabledness has still low comprehension of programming, does not yet understand the syntax of lines of code completely, does not yet know any concepts of iteration, and is not yet able to formulate it generally. However, improving the novice’s understanding requires fundamental knowledge about the meaning of the code or the ability to interact with the algorithm to observe changes in the result. As is generally the case in all phases of an explanation and with regard to all forms of understanding, there are cognitive and motivational prerequisites of understanding at the beginning of an explanation. In the following, we focus in particular on prior knowledge and motivation for the situation of shallow comprehension and shallow enabledness. We would like to emphasize how an explanation comes about. In the situation that is usually at the beginning of a domain-specific explanation, prior knowledge, which is not only domain-specific, plays a crucial role. By looking at prior knowledge in this section in particular, we emphasize the importance of other, more general knowledge to characterize the basic knowledge in this situation. With the increase in understanding in the course of an explanation, does not only the explanandum change (cf. [66])) but also the specific prior knowledge. Motivation is also a prerequisite for the (continuation of the) explanation in all the forms of understanding discussed in this paper. We present them especially in this section to show that motivation is essential to strive for an explanation as motivation is noteworthy for all actions (cf., [47]).
@@
Prior knowledge is thought to influence co-construction between the explainer and the explainee [37] because the explainer relies on the explainee’s currently assumed prior knowledge to scaffold the explainee’s understanding. However, a greater amount of prior knowledge on the part of both the explainer and the explainee does not necessarily lead to better explanations [95].The influence of prior knowledge in explanations seems to be similar to its influence in learning scenarios: prior knowledge is not necessarily beneficial for the explanation process [86]. It was found that the effect of explanation and the amount of prior knowledge on learning are neither independent nor subadditive [95]. Rather, they appear to have a nonlinear influence on learning. An explanation guides the focus in acquiring new knowledge, and prior knowledge helps to identify patterns and constraints. Not all prior knowledge is used as by explainers or explainees, only parts relevant to a task are activated [30]. However, these ‘relevant’ parts are relative to the individual holding the prior knowledge [2]. Thus, the personal relevant knowledge is not necessarily identical for the explainer and the explainee; neither is the assumed prior knowledge of the explainee on the side of the explainer necessarily correct. The assumed knowledge is part of a mental representation that the explainer has of his interaction partner, the partner model [17].
@@
The literature presents a variety of methods for the scientific assessment of prior knowledge. These include verbal methods such as ‘think-aloud’ tasks, verbal protocols and audit trails, and performance metrics on problem solving and troubleshooting or retention over time, as well as methods specific to human-computer interaction, system usage, user explanations of the system, and user prediction of system performance [84]. The simplest method is to query only the domain-specific prior knowledge [12]. However, this method discards most of the total prior knowledge.
An alternative approach is presented by Cole et al. [26], who use a combination of three different questionnaires to assess prior knowledge. The first questionnaire tests prior knowledge using true/false questions about domain-specific conceptual knowledge. In addition, they ask respondents to estimate their prior knowledge, such as “How well do you know X?” The last questionnaire tests prior knowledge by asking about familiarity with the domain, e.g., “How often have you used X?” In addition, the authors provide a questionnaire that asks about attitudes towards the domain, as they influence the explanation [89] as much as the actual knowledge the participants bring to an explanation.
In addition to prior knowledge, the motivation brought to an understanding process also influences its course. Wade and Kidd [91], for example, have shown a bidirectional relationship between curiosity and learning. Therefore, the motivational prerequisites for comprehension will be discussed in the following.
@@
To identify motivational prerequisites, a classical approach describes motivation as influenced by expectancy and value. Expectancy refers to “beliefs about how well they will do at the activity” (Wigfield and Eccles 94, p. 68, cf. Eccles and Wigfield 32). The subjective value of understanding is divided into at least two parts: understanding can have (1) an intrinsic value of the activity of understanding, arising from an interest in the explanandum or a relation to the explainee in the social situation, or (2) understanding the explanation can be useful or valuable for achieving a goal, and thus has an extrinsic value that does not lie in the action itself. For example, understanding digital artifacts is useful or even required in the context of school, university, and the labor market [e.g. 10]. Expectation as well as value depend on the previous experiences of the explainee [14]. With the experience that it is difficult to follow an explanation or that digital artifacts are difficult to understand, one’s self-efficacy to understand XAI-explanations is likely to be low. On the other hand, an adequate explanation increases the self-efficacy to use (digital) technology (Weitz 92; cf. technology acceptance models: Davis 29, Marangunić and Granić 64). Both expectancy and value depend on the explainee’s goals (Eccles and Wigfield 32, cf. Covington 28), such as learning the present progressive tense for a test in school or for living in a foreign country.
@@ 3.2 Shallow comprehension, deep enabledness
Besides the motivational prerequisites already explained (Section 3.1, the access to external resources (e.g., tools such as web search, chatbots, FAQs, other people) may be relevant to be enabled to perform an action without necessarily having a conceptual comprehension of the task. One example is to program another while-loop and make a query to a chatbot (as an external resource) to support this task. In this sense, a deeper enabledness is characterized by flexibility and adaptivity in the use of the external resources (e.g., novices vs. experts, see Section 4.1). The more flexibly an explainee uses the resources and adapts their use to different contextual environments, the deeper is the enabledness. Flexibility here means (i) various resources are known, and (ii) contextual factors (such as time restrictions, current (un)availability of an external resource etc.) are taken into account to choose the appropriate resources in a certain situation. Adaptivity means the ability to select the appropriate resources in novel contexts/in different contextual environments. In this respect, deep enabledness allows someone to use their skills in a wider range of applications, and without the need for resources being present in the immediate environment, because they can be simulated using internal mental representations, e.g., by memorization (see below). However, the deepest form of enabledness cannot be reached without any comprehension, since choosing what the best/most helpful resource is in a certain situation, or making a concrete query when using systems such as web search or chatbots, certainly presupposes some conceptual knowledge of the problem or task (cf. Graesser and Person 42 on the relation between asking deep-level questions and deepening of conceptual knowledge).
@@
In contrast, shallow enabledness encompasses rather basic actions of an individual. This has, for example, been described for early infant development. Mandler [63] proposes that an infant might be able to identify an object as belonging to a category but this cognitive operation is based on perceptual information only and does not require an interpretation of spatiotemporal information. A basic action is embedded in the external representations itself and is externally driven (i.e., there is a perfect match between the physical disposition and the environment; see Identification/recognition in the listed abilities). In contrast to deeper enabledness, only a limited number of external resources are used and they are used context-dependent, in a given way (e.g., using the same query again) and cannot be transferred to another or a changed context. The transition from shallow to a deeper enabledness comes by experience, i.e., a higher degree of flexibility and adaptivity is achieved by richer experience (through repetition in the same or only slightly changed contexts) and a far transfer of skill [36]. Enabledness without deeper comprehension may also be achieved by memorizing a number of steps, procedures or facts (rote learning) without understanding the concepts behind it. However, this approach to enabledness, that Boaler and Greeno [13, p. 9] call a “ritualistic act of knowledge reproduction”, may hinder progress in deepening comprehension because explainees may be less engaged with the underlying concepts. Deeper comprehension, however, requires “thinking practices” [13, p. 9, for mathematical learning]. Therefore, we assume that deeper enabledness as we have defined it, including a flexible use in different contexts, cannot be accomplished by memorization or rote learning.
@@
To define comprehension, among others we made use of the taxonomies proposed by Bloom [59] and Krathwohl [6, 58] to classify learning objectives. They describe different cognitive processes that a learner is able to perform within a learning process. However, originally there is no clear distinction between enabledness and comprehension in these learning taxonomies: Bloom’s taxonomy [59] starts with early processing mechanisms that we ascribe to enabledness but not comprehension, whereas Krathwohl’s taxonomy [6, 58] seems to require interpretation and memory from the beginning. To make the early steps of understanding clear, we will therefore refer to Bloom’s taxonomy. Overall, we recognize that there is a progression from being bound to perceptual information or the context to decontextualized knowledge that is more flexible and adaptive. We view the boundedness to external resources not only to be more shallow in terms of enabledness (see above) but also in terms of comprehension. Therefore, the abilities described in the first two stages on Bloom’s taxonomy seem to belong to shallow comprehension/shallow enabledness (see Section 3.1). It is difficult, however, to make a clear cut between comprehension and enabledness because the different forms of comprehension only become observable by abilities, and therefore by forms of enabledness.
@@
Krathwohl et al. [59] identified the following abilities as reflecting the first steps of comprehension: Repeating, listing, labeling, completing, and selecting. We propose that all of them afford a more or less prepared context (i.e., external resources) in which the knowledge is triggered to be applied. According to Mandler [63], the operations at this level may be based on perceptual information only and do not require an interpretation of the context. For the other abilities, based on the rich context, only little knowledge or comprehension is required to achieve an act (e.g., to label something or to fill something out).
@@
The second level, called "understanding" in Krathwohl’s taxonomy [6], represents, among others, the abilities of interpreting, exemplifying, and summarizing. We propose that they require some knowledge to be connected to another entity that is obvious or present in the environment (i.e., contextualization), thus requiring a deeper comprehension than in the first stage. Contextualization was recognized as a form of interpretation of an activity based on cues from the obvious/present and related to the process of inferencing, which itself depends on the context of its occurrence [8, p. 30]. For our example 1, a reasonable explanation for why the first result is ‘99 bottles …’ is that counter is initialized with ‘99’. This explanation simply accounts for what can be given by referring to the present context, without having to add anything to it. A deeper reason goes beyond the obvious and offers a connection to what cannot be perceived, such as what the output would be if the code in line 1 or 6 were changed in a certain way. Making connections is important in the learning process [38].
@@
All of the three above-mentioned steps may be time-dependent – comparable to cognitive mechanisms of mapping discussed for word-learning in infancy [97], i.e., fast-mapping (short-term; less demanding) and slow-mapping (requiring long-term memory; more demanding), or to different forms of cognitive engagement of students [24, p. 221]. Chi and Wylie [24] differentiate between recall (what we see as shallow comprehension), apply (at the intersection from shallow to deep comprehension), transfer and co-create (deep comprehension).
In the following, we argue that comprehension and enabledness are intertwined. We presume that most types of deep enabledness would build on some kind of comprehension, and, vice versa, deeper enabledness might lead to increased comprehension: (i) comprehension changes depending on enabledness and vice versa, and (ii) comprehension, even shallow, only becomes observable in physical and cognitive abilities that are also signs of enabledness (cf. Section 4.2). Cases of extreme shallow comprehension together with extreme deep enabledness, or cases of shallow comprehension and intermediate enabledness are therefore hard to be defined and would contradict all propositions made above. (iii) There is neurological evidence on language production, according to which comprehension and enabledness cannot be torn apart at the cognitive level because of shared underlying neurological representations, at least from a functional perspective [e.g., 75].
In terms of assessability of comprehension, the question is how verbalization is related to the degree of comprehension. This question arises based on studies on gesture use and learning (cf. extensive work by Goldin-Meadow and colleagues, e.g., Goldin-Meadow 39 for an overview): If comprehension is shallow, a person may not be able to verbalize knowledge, i.e., the knowledge is more implicit. However, Goldin-Meadow and colleagues have shown that other (communicative) behavior such as a person’s gestures can indicate some initial comprehension: even though some learners failed to solve some mathematical problems and were unable to explain their solution, their co-speech gestures conveyed semantic information that was relevant to correctly solve the task but that was not verbalized. Thus, it was assessable nonverbally. The finding that these learners were more likely to benefit from further explanations was interpreted in terms of them having implicit conceptual knowledge. Research on implicit conceptual knowledge has further shown that participants were able to make judgments on whether a problem-solving task is likely to be solvable or not, even when they were not able to provide a correct solution themselves [16]. When comprehension becomes deeper, it becomes explicit, that is, it can be verbalized, and is therefore assessable by asking questions.
@@
The first case is a process at the transition from intermediate to deeper enabledness with alternating more shallow and deeper comprehension (solid line in Figure 2), that may often occur during automation, e.g., in language (L1) acquisition. If something becomes automatized, enabledness becomes deeper, but comprehension may decrease [4, 13], as during the process of automation conceptual knowledge becomes compiled, and therefore less accessible (see Section 2). An example could be automation demonstrated by highly proficient bilinguals in terms of their language switching abilities in everyday life [43]. Research suggests that initially, during the early years of second language acquisition, non-selective language activation occurs. However, as both languages become dominant, the selection rules become automatic. This leads to automatic concept switching, where bilinguals become unaware of the switching rules and rely more on external resources like social cues and a partner’s fluency [43].
@@
The third case (dashed line in Figure 2) is linked to the second one. A progress in enabledness may result in a lagged, nonlinear increase in comprehension following further explanations, that is, the progression in both forms of understanding is stepwise, similar to the idea of bootstrapping from developmental research (Morgan and Demuth 68; cf. Section 4.2). This is based on research on automated processes in second language (L2)learning [cf. 46]. Although L2-learners might encounter difficulties in the process of generating new knowledge, for example the present progressive tense, this process may be supported by establishing relations to previously acquired, related conceptual knowledge, such as the general distinction between present and past, as well as knowledge expressed in identification of (in-)consistent syntactic constructions of the L2 based on their L1 knowledge (cf. Mandler 63; Fiorella and Mayer 35, p. 7). At this stage, comprehension remains behind the intermediate level, but enabledness may increase progressively if L2-learners apply their previous conceptual knowledge by doing and repeating related exercises, e.g., filling missing auxiliaries ‘to be’ or missing -ing suffixes. Learners may also rely on external resources such as textbooks, resources for language practice, or a human tutor to complete the tasks. Learning the different application contexts of the English present progressive tense would enable learners to use the present progressive without recalling basic concepts and to apply it to various similar cases in L2-learning.
@@
Regarding automation, i.e., how efficient an explainee is, different measures could be used, e.g., processing time, or number of errors in relation to total performance. Since we have defined automation as compiled knowledge that may be less accessible to verbalization, the assessment could also consist of two parts, a task and a questionnaire: if the explainee is able to successfully accomplish the task but is unable to explain why or how it was done, comprehension is likely to be shallower. However, this may also be an indicator of transitional conceptual knowledge [40].

### v3

@@ 3.1 Shallow comprehension, shallow enabledness
We see the combination of shallow comprehension and shallow enabledness as a potential starting point for understanding. Thus, this section will focus on the mental state before a change in understanding, or here, before explanation. The first step is to define the underlying conditions and premises of this stage. Any process of understanding should begin with and be tailored to the explainee’s mental representation model, which is built from (prior) own and mediated experiences with the domain and prior knowledge about familiar things. At the beginning of an explanation, explainers and explainees need to establish ’common ground’ (Clark, 2014, p. 109), which involves assessing “what the other person knows at the start, then tailoring one’s own contributions to add to the relevant common ground for that occasion […]”. Common ground is usually based on the prior knowledge that explainees bring to the explanation process. Prior knowledge thus influences the course of the explanation and the understanding that follows: Explainers can build their explanation on it, and explainees’ mental processing of new knowledge is influenced by it (Chi et al., 1994). Therefore, it is particularly interesting to know what prior knowledge is and which parts of it are relevant for an explanation, and how the explainer can identify these relevant parts of prior knowledge. To clarify the situation in this field, we begin with the two examples.
@@
We can imagine that a novice in the field of shallow comprehension and shallow enabledness has still low comprehension of programming, does not yet understand the syntax of lines of code completely, does not yet know any concepts of iteration, and is not yet able to formulate it generally. However, improving the novice’s understanding requires fundamental knowledge about the meaning of the code or the ability to interact with the algorithm to observe changes in the result. As is generally the case in all phases of an explanation and with regard to all forms of understanding, there are cognitive and motivational prerequisites of understanding at the beginning of an explanation. In the following, we focus in particular on prior knowledge and motivation for the situation of shallow comprehension and shallow enabledness. We would like to emphasize how an explanation comes about. In the situation that is usually at the beginning of a domain-specific explanation, prior knowledge, which is not only domain-specific, plays a crucial role. By looking at prior knowledge in this section in particular, we emphasize the importance of other, more general knowledge to characterize the basic knowledge in this situation. With the increase in understanding in the course of an explanation, does not only the explanandum change (cf. (Miller, 2019))) but also the specific prior knowledge. Motivation is also a prerequisite for the (continuation of the) explanation in all the forms of understanding discussed in this paper. We present them especially in this section to show that motivation is essential to strive for an explanation as motivation is noteworthy for all actions (cf., (Heckhausen and Heckhausen, 2018)).
@@
Prior knowledge is thought to influence co-construction between the explainer and the explainee (Gentile et al., 2021) because the explainer relies on the explainee’s currently assumed prior knowledge to scaffold the explainee’s understanding. However, a greater amount of prior knowledge on the part of both the explainer and the explainee does not necessarily lead to better explanations (Williams and Lombrozo, 2010).The influence of prior knowledge in explanations seems to be similar to its influence in learning scenarios: prior knowledge is not necessarily beneficial for the explanation process (Simonsmeier et al., 2021). It was found that the effect of explanation and the amount of prior knowledge on learning are neither independent nor subadditive (Williams and Lombrozo, 2010). Rather, they appear to have a nonlinear influence on learning. An explanation guides the focus in acquiring new knowledge, and prior knowledge helps to identify patterns and constraints. Not all prior knowledge is used as by explainers or explainees, only parts relevant to a task are activated (Dochy and Alexander, 1995). However, these ‘relevant’ parts are relative to the individual holding the prior knowledge (Alba and Hasher, 1983). Thus, the personal relevant knowledge is not necessarily identical for the explainer and the explainee; neither is the assumed prior knowledge of the explainee on the side of the explainer necessarily correct. The assumed knowledge is part of a mental representation that the explainer has of his interaction partner, the partner model (Brennan and Hanna, 2009).
@@
The literature presents a variety of methods for the scientific assessment of prior knowledge. These include verbal methods such as ‘think-aloud’ tasks, verbal protocols and audit trails, and performance metrics on problem solving and troubleshooting or retention over time, as well as methods specific to human-computer interaction, system usage, user explanations of the system, and user prediction of system performance (Sasse, 1991). The simplest method is to query only the domain-specific prior knowledge (Best et al., 2004). However, this method discards most of the total prior knowledge.
An alternative approach is presented by Cole et al. (1986), who use a combination of three different questionnaires to assess prior knowledge. The first questionnaire tests prior knowledge using true/false questions about domain-specific conceptual knowledge. In addition, they ask respondents to estimate their prior knowledge, such as “How well do you know X?” The last questionnaire tests prior knowledge by asking about familiarity with the domain, e.g., “How often have you used X?” In addition, the authors provide a questionnaire that asks about attitudes towards the domain, as they influence the explanation (Tobias, 1994) as much as the actual knowledge the participants bring to an explanation.
In addition to prior knowledge, the motivation brought to an understanding process also influences its course. Wade and Kidd (2019), for example, have shown a bidirectional relationship between curiosity and learning. Therefore, the motivational prerequisites for comprehension will be discussed in the following.
@@
To identify motivational prerequisites, a classical approach describes motivation as influenced by expectancy and value. Expectancy refers to “beliefs about how well they will do at the activity” (Wigfield and Eccles 2000, p. 68, cf. Eccles and Wigfield 2020). The subjective value of understanding is divided into at least two parts: understanding can have (1) an intrinsic value of the activity of understanding, arising from an interest in the explanandum or a relation to the explainee in the social situation, or (2) understanding the explanation can be useful or valuable for achieving a goal, and thus has an extrinsic value that does not lie in the action itself. For example, understanding digital artifacts is useful or even required in the context of school, university, and the labor market (e.g. Ball et al., 2017). Expectation as well as value depend on the previous experiences of the explainee (Bonanati and Buhl, 2021). With the experience that it is difficult to follow an explanation or that digital artifacts are difficult to understand, one’s self-efficacy to understand XAI-explanations is likely to be low. On the other hand, an adequate explanation increases the self-efficacy to use (digital) technology (Weitz 2021; cf. technology acceptance models: Davis 1989, Marangunić and Granić 2014). Both expectancy and value depend on the explainee’s goals (Eccles and Wigfield 2020, cf. Covington 2000), such as learning the present progressive tense for a test in school or for living in a foreign country.
@@ 3.2 Shallow comprehension, deep enabledness
Besides the motivational prerequisites already explained in Section 3.1, the access to external resources (e.g., tools such as web search, chatbots, FAQs, other people) may be relevant to be enabled to perform an action without necessarily having a conceptual comprehension of the task. One example is to program another while-loop and make a query to a chatbot (as an external resource) to support this task. In this sense, a deeper enabledness is characterized by flexibility and adaptivity in the use of the external resources (e.g., novices vs. experts, see Section 4.1). The more flexibly an explainee uses the resources and adapts their use to different contextual environments, the deeper is the enabledness. Flexibility here means (i) various resources are known, and (ii) contextual factors (such as time restrictions, current (un)availability of an external resource etc.) are taken into account to choose the appropriate resources in a certain situation. Adaptivity means the ability to select the appropriate resources in novel contexts/in different contextual environments. In this respect, deep enabledness allows someone to use their skills in a wider range of applications, and without the need for resources being present in the immediate environment, because they can be simulated using internal mental representations, e.g., by memorization (see below). However, the deepest form of enabledness cannot be reached without any comprehension, since choosing what the best/most helpful resource is in a certain situation, or making a concrete query when using systems such as web search or chatbots, certainly presupposes some conceptual knowledge of the problem or task (cf. Graesser and Person 1994 on the relation between asking deep-level questions and deepening of conceptual knowledge).
@@
In contrast, shallow enabledness encompasses rather basic actions of an individual. This has, for example, been described for early infant development. Mandler (2012) proposes that an infant might be able to identify an object as belonging to a category but this cognitive operation is based on perceptual information only and does not require an interpretation of spatiotemporal information. A basic action is embedded in the external representations itself and is externally driven (i.e., there is a perfect match between the physical disposition and the environment; see Identification/recognition in the listed abilities). In contrast to deeper enabledness, only a limited number of external resources are used and they are used context-dependent, in a given way (e.g., using the same query again) and cannot be transferred to another or a changed context. The transition from shallow to a deeper enabledness comes by experience, i.e., a higher degree of flexibility and adaptivity is achieved by richer experience (through repetition in the same or only slightly changed contexts) and a far transfer of skill (Gardner, 1999). Enabledness without deeper comprehension may also be achieved by memorizing a number of steps, procedures or facts (rote learning) without understanding the concepts behind it. However, this approach to enabledness, that Boaler and Greeno (2000, p. 9) call a “ritualistic act of knowledge reproduction”, may hinder progress in deepening comprehension because explainees may be less engaged with the underlying concepts. Deeper comprehension, however, requires “thinking practices” (Boaler and Greeno, 2000, p. 9, for mathematical learning). Therefore, we assume that deeper enabledness as we have defined it, including a flexible use in different contexts, cannot be accomplished by memorization or rote learning.
@@
To define comprehension, among others we made use of the taxonomies proposed by Bloom (Krathwohl et al., 1969) and Krathwohl (Anderson and Krathwohl, 2001; Krathwohl, 2002) to classify learning objectives. They describe different cognitive processes that a learner is able to perform within a learning process. However, originally there is no clear distinction between enabledness and comprehension in these learning taxonomies: Bloom’s taxonomy (Krathwohl et al., 1969) starts with early processing mechanisms that we ascribe to enabledness but not comprehension, whereas Krathwohl’s taxonomy (Anderson and Krathwohl, 2001; Krathwohl, 2002) seems to require interpretation and memory from the beginning. To make the early steps of understanding clear, we will therefore refer to Bloom’s taxonomy. Overall, we recognize that there is a progression from being bound to perceptual information or the context to decontextualized knowledge that is more flexible and adaptive. We view the boundedness to external resources not only to be more shallow in terms of enabledness (see above) but also in terms of comprehension. Therefore, the abilities described in the first two stages on Bloom’s taxonomy seem to belong to shallow comprehension/shallow enabledness (see Section 3.1). It is difficult, however, to make a clear cut between comprehension and enabledness because the different forms of comprehension only become observable by abilities, and therefore by forms of enabledness.
@@
Krathwohl et al. (1969) identified the following abilities as reflecting the first steps of comprehension: Repeating, listing, labeling, completing, and selecting. We propose that all of them afford a more or less prepared context (i.e., external resources) in which the knowledge is triggered to be applied. According to Mandler (2012), the operations at this level may be based on perceptual information only and do not require an interpretation of the context. For the other abilities, based on the rich context, only little knowledge or comprehension is required to achieve an act (e.g., to label something or to fill something out).
@@
The second level, called ”understanding” in Krathwohl’s taxonomy (Anderson and Krathwohl, 2001), represents, among others, the abilities of interpreting, exemplifying, and summarizing. We propose that they require some knowledge to be connected to another entity that is obvious or present in the environment (i.e., contextualization), thus requiring a deeper comprehension than in the first stage. Contextualization was recognized as a form of interpretation of an activity based on cues from the obvious/present and related to the process of inferencing, which itself depends on the context of its occurrence (Auer, 1992, p. 30). For our example 1, a reasonable explanation for why the first result is ‘99 bottles …’ is that counter is initialized with ‘99’. This explanation simply accounts for what can be given by referring to the present context, without having to add anything to it. A deeper reason goes beyond the obvious and offers a connection to what cannot be perceived, such as what the output would be if the code in line 1 or 6 were changed in a certain way. Making connections is important in the learning process (Gentner et al., 2003).
@@
All of the three above-mentioned steps may be time-dependent – comparable to cognitive mechanisms of mapping discussed for word-learning in infancy (Wojcik, 2013), i.e., fast-mapping (short-term; less demanding) and slow-mapping (requiring long-term memory; more demanding), or to different forms of cognitive engagement of students (Chi and Wylie, 2014, p. 221). Chi and Wylie (2014) differentiate between recall (what we see as shallow comprehension), apply (at the intersection from shallow to deep comprehension), transfer and co-create (deep comprehension).
In the following, we argue that comprehension and enabledness are intertwined. We presume that most types of deep enabledness would build on some kind of comprehension, and, vice versa, deeper enabledness might lead to increased comprehension: (i) comprehension changes depending on enabledness and vice versa, and (ii) comprehension, even shallow, only becomes observable in physical and cognitive abilities that are also signs of enabledness (cf. Section 4.2). Cases of extreme shallow comprehension together with extreme deep enabledness, or cases of shallow comprehension and intermediate enabledness are therefore hard to be defined and would contradict all propositions made above. (iii) There is neurological evidence on language production, according to which comprehension and enabledness cannot be torn apart at the cognitive level because of shared underlying neurological representations, at least from a functional perspective (e.g., Pickering and Garrod, 2013).
In terms of assessability of comprehension, the question is how verbalization is related to the degree of comprehension. This question arises based on studies on gesture use and learning (cf. extensive work by Goldin-Meadow and colleagues, e.g., Goldin-Meadow 2017 for an overview): If comprehension is shallow, a person may not be able to verbalize knowledge, i.e., the knowledge is more implicit. However, Goldin-Meadow and colleagues have shown that other (communicative) behavior such as a person’s gestures can indicate some initial comprehension: even though some learners failed to solve some mathematical problems and were unable to explain their solution, their co-speech gestures conveyed semantic information that was relevant to correctly solve the task but that was not verbalized. Thus, it was assessable nonverbally. The finding that these learners were more likely to benefit from further explanations was interpreted in terms of them having implicit conceptual knowledge. Research on implicit conceptual knowledge has further shown that participants were able to make judgments on whether a problem-solving task is likely to be solvable or not, even when they were not able to provide a correct solution themselves (Bowers et al., 1990). When comprehension becomes deeper, it becomes explicit, that is, it can be verbalized, and is therefore assessable by asking questions.
@@
The first case is a process at the transition from intermediate to deeper enabledness with alternating more shallow and deeper comprehension (solid line in Figure 2), that may often occur during automation, e.g., in language (L1) acquisition. If something becomes automatized, enabledness becomes deeper, but comprehension may decrease (Anderson, 1996; Boaler and Greeno, 2000), as during the process of automation conceptual knowledge becomes compiled, and therefore less accessible (see Section 2). An example could be automation demonstrated by highly proficient bilinguals in terms of their language switching abilities in everyday life (Green and Abutalebi, 2013). Research suggests that initially, during the early years of second language acquisition, non-selective language activation occurs. However, as both languages become dominant, the selection rules become automatic. This leads to automatic concept switching, where bilinguals become unaware of the switching rules and rely more on external resources like social cues and a partner’s fluency (Green and Abutalebi, 2013).
@@
The third case (dashed line in Figure 2) is linked to the second one. A progress in enabledness may result in a lagged, nonlinear increase in comprehension following further explanations, that is, the progression in both forms of understanding is stepwise, similar to the idea of bootstrapping from developmental research (Morgan and Demuth 1996; cf. Section 4.2). This is based on research on automated processes in second language (L2)learning (cf. Hao and Othman, 2021). Although L2-learners might encounter difficulties in the process of generating new knowledge, for example the present progressive tense, this process may be supported by establishing relations to previously acquired, related conceptual knowledge, such as the general distinction between present and past, as well as knowledge expressed in identification of (in-)consistent syntactic constructions of the L2 based on their L1 knowledge (cf. Mandler 2012; Fiorella and Mayer 2015, p. 7). At this stage, comprehension remains behind the intermediate level, but enabledness may increase progressively if L2-learners apply their previous conceptual knowledge by doing and repeating related exercises, e.g., filling missing auxiliaries ‘to be’ or missing -ing suffixes. Learners may also rely on external resources such as textbooks, resources for language practice, or a human tutor to complete the tasks. Learning the different application contexts of the English present progressive tense would enable learners to use the present progressive without recalling basic concepts and to apply it to various similar cases in L2-learning.
@@
Regarding automation, i.e., how efficient an explainee is, different measures could be used, e.g., processing time, or number of errors in relation to total performance. Since we have defined automation as compiled knowledge that may be less accessible to verbalization, the assessment could also consist of two parts, a task and a questionnaire: if the explainee is able to successfully accomplish the task but is unable to explain why or how it was done, comprehension is likely to be shallower. However, this may also be an indicator of transitional conceptual knowledge (Goldin-Meadow et al., 1993).